OpenStudy (anonymous):

Write an equation for the graph of a linear function that: Passes through H(0,-3) and has slope 7/16 ... help please?

OpenStudy (anonymous):

Lets use the slope-point form of a line, which in general terms is written as\[y - y_1 = m(x-x_1)\]where m is the slope and y_1 and x_1 are the coordinates of the point the line passes through. I'll let you plug in the numbers.

OpenStudy (saifoo.khan):

\[y = \frac{7}{16}x-3\]